When the necessary conditions are met, a two-tailed test is being conducted to test the difference between two population proportions. The two sample proportions are p1hat=0.052 and p2hat=0.023 and the standard error of the sampling distribution of p1-p2 is 0.0085. The calculated value of the test statistic will be z = 3.41.

A. True

B. False
Answer #1

Result:

When the necessary conditions are met, a two-tailed test is being conducted to test the difference between two population proportions. The two sample proportions are p1hat=0.052 and p2hat=0.023 and the standard error of the sampling distribution of p1-p2 is 0.0085. The calculated value of the test statistic will be z = 3.41.

Correct option : A. True

B. False

test statistic z = (p1-p2)/se = (0.052-0.023)/0.0085

=3.411764706